# Training Nonlinear Transformers for Chain-of-Thought Inference: A Theoretical Generalization Analysis

### v2

@@ 3.4 CoT generalization guarantee
which measures the minimum probability of the most probable $K$-step reasoning trajectory over the initial TSR pattern. We also define the min-max input-label transition probability as
@@
which measures the minimum probability of the most probable output over the initial TSR pattern.
@@
Given a trained model that satisfies conditions (i) to (iii) in Theorem 1, as long as (iv)
${\bm{\mu}}_{j}^{\prime}\in\text{span}({\bm{\mu}}_{1},{\bm{\mu}}_{2},\cdots,{\bm{\mu}}_{M}),$ (18)
for $j\in[M^{\prime}]$, and (v) the number of testing examples for every task $f\in\mathcal{T}^{\prime}$ is
@@
Theorem 2 characterizes the sufficient conditions for a trained one-layer Transformer to generate all $K$-steps reasoning correctly by CoT for a task $f$ in $\mathcal{T}^{\prime}$. First, the TSR patterns of a new task in $\mathcal{T^{\prime}}$ should be linear combinations of TRR patterns in the training tasks in $\mathcal{T}$. Second, the number of context examples should be in the order of ${\alpha^{\prime}}^{-2}$, ${\rho_{s}^{f}}^{-2}$, and ${\tau^{f}}^{-2}$. One can equivalently interpret the decrease in the number of required context examples to achieve zero CoT error as an improvement of the CoT accuracy with fixed context length. Then, when the fraction $\alpha^{\prime}$ of contexts where the TSR pattern of the first step input is the same as the query increases, the contexts become more informative for the query. Thus, the CoT accuracy increases. When $\rho^{f}$ and $\tau^{f}$ increase, the reasoning labels in the context examples are more likely to be accurate based on their definitions in (11) and (16), then the CoT accuracy is improved.
@@ 3.5 ICL Generalization and Comparison with CoT
Because only input-label pairs are used as context examples without intermediate reasoning steps for ICL, then the input-label pairs in context examples should be accurate on average. Otherwise, the context examples are not informative about the task and will lead to the failure of ICL. We formulate this requirement as Condition 1.
@@
Condition 1 requires that in a context example, if the input TSR is ${\bm{\mu}}_{i}^{\prime}$, then the output TSR needs to be $f({\bm{\mu}}_{i}^{\prime})$ with the largest probability over all other TSR patterns. It is intuitive that the success of ICL requires this condition. Note that although (11) indicates that, $A_{k}^{f}(i,j^{*})$ achieves the largest value for all $j$ when ${\bm{\mu}}_{j^{*}}^{\prime}=f_{k}({\bm{\mu}}_{i}^{\prime})$ for every $k$ and $i$, (11) does not always lead to (20). Example 1 demonstrates a case where Condition 1 does not hold. Given the input ${\bm{\mu}}_{1}^{\prime}$, the correct inference output shall be ${\bm{\mu}}_{2}^{\prime}$, but ${\bm{B}}^{f}$ indicates that the most probable output is ${\bm{\mu}}_{1}^{\prime}$.
@@
Given a trained model that satisfies conditions (i) to (iii) of Theorem 1 and (18), for the testing task $f\in\mathcal{T}^{\prime}$,
  1. a.
    if Condition 1 does not hold, then $\bar{R}_{ICL,{\bm{x}}_{query}\in\mathcal{M}^{\prime},f\in\mathcal{T}^{\prime}}^{f}(\Psi)\geq\Omega(1)$;
  2. b.
    if Condition 1 holds, we have $\bar{R}_{ICL,{\bm{x}}_{query}\in\mathcal{M}^{\prime},f\in\mathcal{T}^{\prime}}^{f}(\Psi)=0$, as long as the number of testing examples is
@@
Ref. Li et al. (2023d) also shows the advantage of CoT over ICL to learn MLP functions, but in a different setting from ours, where our studied tasks operate on patterns. More importantly, this paper characterizes the CoT and ICL performance theoretically when the testing task has a distribution shift from training tasks (TRR patterns to TSR patterns), and the testing examples contain errors, while Li et al. (2023d) only empirically evaluates the CoT and ICL performance with noisy examples.
@@ 4.1 Transformers implement CoT by Attending to the Most Similar Examples Every Step
In this section, we characterize the key mechanism of a properly trained one-layer Transformer to implement CoT on a $K$-steps reasoning task via training dynamics analysis of the attention layer, as demonstrated in Figure 2. This is different from the mechanism study in (Li et al., 2023d; Feng et al., 2023) by constructing a model that can conduct CoT. We have the following proposition for the trained model.
@@ 4.2 An Overview of the Proof
The technical challenges of the proof are concentrated on Theorem 1, where the property of the trained model is derived. The proof of Theorem 1 is built upon three Lemmas, which characterize the two stages of the training dynamics. Specifically, Lemmas 3 and 4 show that if a training prompt ${\bm{P}}$ includes the first $k$ steps of the reasoning query, then the attention weights on columns of ${\bm{P}}$ with a different step from the query decrease to be close to zero in the first stage. Lemma 5 computes the gradient updates in the second stage, where the attention weights on columns in ${\bm{P}}$ that correspond to step $k$ and have the same TRR pattern as the query gradually become dominant. Theorem 1 unveils this training process by showing the required number of training iterations and sample complexity.
To prove Theorem 2, we first compute the required number of context examples for the new task $f\in\mathcal{T}^{\prime}$ so that by concentration inequalities, the number of context examples with accurate TSR is larger than examples with inaccurate TSR patterns in all $K$ reasoning steps with high probability. Then, due to the linear correlation between TSR and TRR patterns (18), we also show that the trained Transformer can attend to context columns with the same TSR pattern as the query. Therefore, the model can make the correct generation in each step. Theorem 3 follows a similar proof idea to Theorem 2, with the difference that the trained model predicts output directly from the input query following ${\bm{B}}^{f}$ instead of using $K$ reasoning steps following ${\bm{A}}_{k}^{f},k\in[K]$ in CoT. Therefore, Condition 1 is required for the success of ICL generalization.
@@ 6 Conclusion, Limitations, and Future Works
Appendix A Additional Discussions
A.1 The motivation to study one-layer single-head Transformers
@@
Second, the state-of-the-art theoretical works (Li et al., 2023a, 2024a; Huang et al., 2023; Makkuva et al., 2024a; Ildiz et al., 2024) on optimization and generalization also focus mainly on one-layer Transformers. No existing works study the optimization and generalization of CoT even for one-layer Transformers. Therefore, we plan to focus on the one-layer analysis to obtain more theoretical insights. We leave the theoretical analysis of the multi-layer case as future works.
@@
A.2 The motivation of the data and task formulation
@@
A.3 The discussion of positional encoding
@@ Appendix C Preliminaries
Appendix D Proof of Main Theorems
D.1 Proof of Theorem 1
@@
$p_{n}(t+1)=\frac{e^{u}}{e^{u}+\frac{1-\alpha}{\alpha}}\gtrsim 1-\frac{1-\alpha}{\alpha}e^{-u},$ (47)
@@
$1-p_{n}(t+1)\geq\frac{1-\alpha}{\alpha e^{u}+(1-\alpha)}\gtrsim\frac{1-\alpha}{\alpha}e^{-u}.$ (48)
Then, we prove that when $t$ is large enough, $u(t)\geq\frac{1}{2}\log\frac{\eta(1-\alpha)^{2}t}{\alpha^{2}M}$. We show it by induction. Suppose that the conclusion holds when $t=t_{0}$, then
$u(t)\geq$ $\frac{\eta}{KM}\sum_{b=0}^{t_{0}}(1-p_{n}(b))^{2}p_{n}(b)+\frac{\eta}{KM}(1-p_{n}(t))^{2}p_{n}(t)$ (49)
$\geq$ $\frac{1}{2}\log\frac{(1-\alpha)^{2}t}{2\alpha^{2}KM}+\frac{\eta}{KM}(1-p_{n}(t))^{2}p_{n}(t)$
$\geq$ $\frac{1}{2}\log\frac{\eta(1-\alpha)^{2}(t+1)}{\alpha^{2}KM},$
@@
$\frac{1}{2}\log(1+\frac{1}{t})\leq\frac{1}{2t}\leq\frac{\eta}{KM}\cdot(\frac{1-\alpha}{\alpha})^{2}e^{-\log\frac{\eta(1-\alpha)^{2}t}{\alpha^{2}KM}}.$ (50)
@@
$(\frac{1-\alpha}{\alpha})^{2}e^{-2u}\leq\epsilon.$ (51)
@@
$u\geq\frac{1}{2}\log\frac{1}{\epsilon}+\log\frac{1-\alpha}{\alpha}.$ (52)
@@
$\frac{1}{2}\log\frac{\eta t}{KM}+\log\frac{1-\alpha}{\alpha}\geq\frac{1}{2}\log\frac{1}{\epsilon}+\log\frac{1-\alpha}{\alpha},$ (53)
@@
${\tilde{{\bm{p}}_{i}^{n}}}^{\top}{\bm{W}}^{(t)}\tilde{{\bm{p}}}_{query}^{n}\geq\log\frac{K}{\epsilon}.$ (55)
@@
${\tilde{{\bm{p}}_{i}^{n}}}^{\top}{\bm{W}}^{(t)}\tilde{{\bm{p}}}_{query}^{n}\geq\Theta(1)\cdot\log\frac{K}{\epsilon}=\Theta(\log\frac{K}{\epsilon}).$ (56)
@@ D.2 Proof of Theorem 2
$\tilde{{\bm{p}}_{i}}=\begin{pmatrix}{\bm{\mu}}_{j}^{\prime}\\ {\bm{y}}_{i}\\ \end{pmatrix}+\begin{pmatrix}{\bm{\delta}}_{i}\\ \bm{0}\\ \end{pmatrix}+{\bm{c}}_{i}.$ (65)
@@
$=$ $(({{\bm{\mu}}_{j}^{\prime}}^{\top},{\bm{y}}_{i}^{\top})+{\bm{c}}_{i}^{\top}))({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})\tilde{{\bm{p}}}_{query}.$
Let ${\bm{\mu}}_{j}^{\prime}=\sum_{i=1}^{M^{\prime}}\lambda_{j,i}{\bm{\mu}}_{i}$. Then, we have
@@
$=$ $((\sum_{i=1}^{M^{\prime}}\lambda_{j,i}{\bm{\mu}}_{i}^{\top},{\bm{y}}_{i}^{\top})+{\bm{c}}_{i}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})((\sum_{i=1}^{M^{\prime}}\lambda_{j,i}{\bm{\mu}}_{i}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$
$=$ $\sum_{i=1}^{M^{\prime}}\lambda_{j,i}^{2}(({\bm{\mu}}_{i}^{\top},{\bm{y}}_{i}^{\top})+{\bm{c}}_{i}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})(({\bm{\mu}}_{i}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$
$+\sum_{i\neq i^{\prime}}\lambda_{j,i}\lambda_{j,i^{\prime}}({\bm{\mu}}_{i}^{\top},{\bm{y}}_{i}^{\top},{\bm{c}}_{i}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})(({\bm{\mu}}_{i^{\prime}}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$
$\geq$ $\Theta(\log\frac{K}{\epsilon})-\epsilon$
$=$ $\Theta(\log\frac{K}{\epsilon}),$
where the second to last step is by Theorem 1. Since the gradient updates for different TRR patterns are very close to each other, we have that if $\sum_{i=1}^{M^{\prime}}\lambda_{j,i}\lambda_{k,i}=0$,
$\sum_{i=1}^{M^{\prime}}\lambda_{j,i}\lambda_{j,i^{\prime}}(({\bm{\mu}}_{i}^{\top},{\bm{y}}_{i}^{\top})+{\bm{c}}_{i}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})(({\bm{\mu}}_{i}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$ (70)
$\lesssim$ $\epsilon\log\frac{K}{\epsilon}.$
Hence, for $\tilde{{\bm{p}}_{i}}$ that shares a different TSR pattern with $\tilde{{\bm{p}}_{i}}$,
$\tilde{{\bm{p}}_{i}}^{\top}{\bm{W}}^{(T)}\tilde{{\bm{p}}}_{query}\lesssim\epsilon\log\frac{K}{\epsilon}.$ (71)
@@
$\tilde{{\bm{p}}}_{j}^{\top}{\bm{W}}^{(T)}(({\bm{v}}_{k}^{\top},\bm{0}^{\top})^{\top}+{\bm{c}}_{k}^{\top})^{\top}=\Theta(\log\frac{K}{\epsilon}),$ (74)

### v3

@@ 3.4 CoT generalization guarantee
which measures the minimum probability, over all the initial TSR patterns, of the $K$-step reasoning trajectory that has the highest probability over all $K$-step trajectories. We also define the min-max input-label transition probability as
@@
which measures the minimum probability, over all the initial TSR patterns, of the output that has the highest probability over outputs.
@@
Given a trained model, the training process of which satisfies conditions (i) to (iii) in Theorem 1, then as long as
(iv) each TSR pattern ${\bm{\mu}}_{j}^{\prime}$ in the orthonormal set $\{{\bm{\mu}}_{j}^{\prime}\}_{j=1}^{M^{\prime}}$ satisfies
${\bm{\mu}}_{j}^{\prime}=\bm{\lambda}_{j}+\tilde{{\bm{\mu}}}_{j}$ (18)
where $\bm{\lambda}_{j}\perp\text{span}({\bm{\mu}}_{1},\cdots,{\bm{\mu}}_{M})$, $\tilde{{\bm{\mu}}}_{j}\in\text{span}({\bm{\mu}}_{1},\cdots,{\bm{\mu}}_{M})$, and $\|\tilde{{\bm{\mu}}}_{j}\|\geq\Theta((\log\epsilon^{-1})^{-1})$,
and (v) the number of testing examples for any $f\in\mathcal{T}^{\prime}$ is
@@
Theorem 2 proves that a trained one-layer Transformer can generate all $K$-steps reasoning correctly by CoT for a new task $f$ in $\mathcal{T}^{\prime}$ with two additional conditions. Condition (iv) means that each TSR pattern in the task set $\mathcal{T^{\prime}}$ is the summation of a component that belongs to the span of the TRR patterns and a component that is perpendicular to the span.
Condition (v) indicates that, to achieve the desired CoT accuracy, the number of context examples should be proportional to ${\alpha^{\prime}}^{-2}$, ${\rho_{s}^{f}}^{-2}$, and ${\tau^{f}}^{-2}$, meaning it decreases as $\alpha^{\prime}$, $\rho_{s}^{f}$, or $\tau_{s}^{f}$ increase. It can be interpreted as follows, if the number of context examples remains fixed, an increase in $\alpha^{\prime}$, $\rho_{s}^{f}$, or $\tau_{s}^{f}$ results in improved CoT accuracy. This aligns with intuition, because $\alpha^{\prime}$ represents the fraction of examples similar to the query, and $\rho^{f}$ and $\tau^{f}$ reflect the accuracy of the reasoning steps in the context examples.
@@ 3.5 ICL Generalization and Comparison with CoT
Because only input-label pairs are used as context examples for ICL, the input-label pairs in context examples should be accurate overall to be informative about the task. We formulate this requirement as Condition 1.
@@
Condition 1 requires that in a context example, if the input TSR is ${\bm{\mu}}_{i}^{\prime}$, then $f({\bm{\mu}}_{i}^{\prime})$ is the output TSR pattern with the highest probability over all TSR patterns. Note that (11) indicates that, for every $k$ and $i$, when ${\bm{\mu}}_{i}^{\prime}$ is the $k$-th step input, $f_{k}({\bm{\mu}}_{i}^{\prime})$ is the step-$k$ output with the highest probability over all TSR patterns. However, (11) does not necessarily imply (20). In Example 1, given the input ${\bm{\mu}}_{1}^{\prime}$, although the inference trajectory ${\bm{\mu}}_{1}^{\prime}\rightarrow{\bm{\mu}}_{1}^{\prime}\rightarrow{\bm{\mu}}_{2}^{\prime}$ under $f$ has the highest probability over all $2$-step trajectories, ${\bm{\mu}}_{1}^{\prime}$ has the higher probability to be the final output than the correct output ${\bm{\mu}}_{2}^{\prime}$ by the two-step transition matrix ${\bm{B}}^{f}$, thus violating Condition 1.
@@
Given a trained model, the training process of which satisfies conditions (i) to (iii) of Theorem 1 and (18), for the testing task $f\in\mathcal{T}^{\prime}$,
  1. Case A.
    if Condition 1 does not hold, then $\bar{R}_{ICL,{\bm{x}}_{query}\in\mathcal{M}^{\prime},f\in\mathcal{T}^{\prime}}^{f}(\Psi)\geq\Omega(1)$, no matter how large the number of training samples $l_{ts}^{f}$ is;
  2. Case B.
    if Condition 1 holds, then $\bar{R}_{ICL,{\bm{x}}_{query}\in\mathcal{M}^{\prime},f\in\mathcal{T}^{\prime}}^{f}(\Psi)=0$, provided that
@@
Ref. Li et al. (2023c) also shows the advantage of CoT over ICL to learn MLP functions, but in a different setting from ours, where our studied tasks operate on patterns. More importantly, this paper characterizes the CoT and ICL performance theoretically when the testing task has a distribution shift from training tasks (TRR patterns to TSR patterns), and the testing examples contain errors, while Li et al. (2023c) only empirically evaluates the CoT and ICL performance with noisy examples.
@@ 4.1 Transformers implement CoT by Attending to the Most Similar Examples Every Step
Figure 2: Concentration of attention weights for CoT inference.
We characterize the key mechanism of a properly trained one-layer Transformer to implement CoT on a $K$-steps reasoning task via training dynamics analysis of the attention layer, as demonstrated in Figure 2. This is different from the mechanism study in (Li et al., 2023c; Feng et al., 2023) by constructing a model that can conduct CoT. We have the following proposition for the trained model.
@@ 4.2 An Overview of the Proof
The technical challenges of the proof are concentrated on Theorem 1, where the property of the trained model is derived. The proof of Theorem 1 is built upon three Lemmas, which characterize the two stages of the training dynamics, i.e., Transformers first attend to tokens with the same step as the query and then, among them, further concentrate on tokens that share the same TSR pattern as the query. Specifically, Lemmas 3 and 4 show that if a training prompt ${\bm{P}}$ includes the first $k$ steps of the reasoning query, then the attention weights on columns of ${\bm{P}}$ with a different step from the query decrease to be close to zero in the first stage. Lemma 5 computes the gradient updates in the second stage, where the attention weights on columns in ${\bm{P}}$ that correspond to the same step and have the same TRR pattern as the query gradually become dominant. Theorem 1 unveils this training process by showing the required number of training iterations and sample complexity.
To prove Theorem 2, we first compute the required number of context examples for the new task $f\in\mathcal{T}^{\prime}$ so that by concentration inequalities, the number of context examples with accurate TSR is larger than examples with inaccurate TSR patterns in all $K$ reasoning steps with high probability. Then, by the correlation between TRR and TSR patterns (18), we also show that the trained Transformer can attend to context columns with the same TSR pattern as the query. Therefore, the model can make the correct generation in each step. Theorem 3 follows a similar proof idea to Theorem 2, with the difference that the trained model predicts output directly from the input query following ${\bm{B}}^{f}$ instead of ${\bm{A}}_{k}^{f},k\in[K]$ in CoT. Therefore, Condition 1 is required for the success of ICL generalization.
@@ 6 Conclusion, Limitations, and Future Works
Acknowledgments
This work was supported by National Science Foundation(NSF) #2430223, Army Research Office (ARO) W911NF-25-1-0020, and the Rensselaer-IBM Future of Computing Research Collaboration (http://airc.rpi.edu). We also thank all anonymous reviewers for their constructive comments.
@@ Acknowledgments
Appendix A Experiments on Real-World Data
We consider a simple arithmetic task that outputs $((A_{1}o_{1}A_{2})o_{2}A_{3})o_{3}A_{4}$ given $A_{1},A_{2},A_{3},A_{4}$ chosen from integers from 00 to $9$ as the input, where $o_{1},o_{2},o_{3}\in O=\{+,-,\times\}$. The CoT output follows the format of $A_{1}o_{1}A_{2}=S_{1},S_{1}o_{2}A_{3}=S_{2},S_{2}o_{3}A_{4}=S_{3}$ and will be evaluated by whether all the three steps are correct for the query as (7). ICL directly outputs $S_{3}$, and the performance is evaluated by the prediction accuracy of $S_{3}$ as (9). In the following experimental settings, the accuracy is computed on $50$ prompts. Each prompt contains three context examples. The inference model is GPT-4 (OpenAI, 2023).
An increasing number of erroneous examples hurts the CoT generalization. To model the errors in the context examples in the testing prompt, we replace $o_{3}$ with one operation $\hat{o_{3}}$ from $O\backslash o_{3}$ in the presentation of some of the context examples in the testing prompt. Note that the output values $S_{3}$ are still correctly computed from $S_{3}=S_{2}o_{3}A_{4}$. Table 1 shows that when the total number of testing examples is fixed to be three, with the increasing number of incorrect examples, the testing accuracy decreases. This is consistent with Remark 1 for Theorem 2.
##\## of incorrect examples 0 1 2 3
CoT accuracy $100\%$ $100\%$ $56\%$ $0\%$
Table 1: The accuracy with different numbers of incorrect examples for CoT. Errors in presenting $o_{3}$.
CoT is more robust to erroneous examples with implementation error than ICL. In this setting, the error in a context examples is introduced by replacing $o_{1}$ with one operation $\hat{o}_{1}$ randomly and independently selected from $O\backslash o_{1}$. Hence, $S_{1}=A_{1}\hat{o}_{1}A_{2}$, and the successive computation are based on the wrongly computed $S_{1}$. The results in Table 2 shows that when two incorrect examples exist, CoT performs better than ICL, which justifies Remark 2 for Theorem 3.
##\## of incorrect examples 0 1 2
CoT accuracy $100\%$ $100\%$ $100\%$
ICL accuracy $100\%$ $100\%$ $60\%$
Table 2: The accuracy with different numbers of incorrect examples for CoT and ICL. Errors in implementing $o_{1}$.
Appendix B Additional Discussions
B.1 The motivation to study one-layer single-head Transformers
@@
Second, the state-of-the-art theoretical works (Li et al., 2023a; 2024a; Huang et al., 2023; Makkuva et al., 2024a; Ildiz et al., 2024) on optimization and generalization also focus mainly on one-layer Transformers. No existing works study the optimization and generalization of CoT even for one-layer Transformers. Therefore, we plan to focus on the one-layer analysis to obtain more theoretical insights. We leave the theoretical analysis of the multi-layer case as future works.
@@
B.2 The motivation of the data and task formulation
@@
B.3 The discussion of positional encoding
@@ Appendix D Preliminaries
Appendix E Proof of Main Theorems
E.1 Proof of Theorem 1
@@
$p_{n}(t+1)=\frac{e^{u}}{e^{u}+\frac{1-\frac{\alpha}{K}}{\frac{\alpha}{K}}}\gtrsim 1-\frac{1-\frac{\alpha}{K}}{\frac{\alpha}{K}}e^{-u},$ (47)
@@
$1-p_{n}(t+1)\geq\frac{1-\frac{\alpha}{K}}{\frac{\alpha}{K}e^{u}+(1-\frac{\alpha}{K})}\gtrsim\frac{1-\frac{\alpha}{K}}{\frac{\alpha}{K}}e^{-u}.$ (48)
Then, we prove that when $t$ is large enough, $u(t)\geq\frac{1}{2}\log\frac{\eta(1-\alpha)^{2}t}{\alpha^{2}KM}$. We show it by induction. Suppose that the conclusion holds when $t=t_{0}$, then
$u(t+1)\geq$ $\frac{\eta}{KM}\sum_{b=0}^{t_{0}}(1-p_{n}(b))^{2}p_{n}(b)+\frac{\eta}{KM}(1-p_{n}(t))^{2}p_{n}(t)$ (49)
$\geq$ $\frac{1}{2}\log\frac{(K-\alpha)^{2}t}{2\alpha^{2}KM}+\frac{\eta}{KM}(1-p_{n}(t))^{2}p_{n}(t)$
$\geq$ $\frac{1}{2}\log\frac{\eta(K-\alpha)^{2}(t+1)}{\alpha^{2}KM},$
@@
$\frac{1}{2}\log(1+\frac{1}{t})\leq\frac{1}{2t}\leq\frac{\eta}{KM}\cdot(\frac{K-\alpha}{\alpha})^{2}e^{-\log\frac{\eta(K-\alpha)^{2}t}{\alpha^{2}KM}}.$ (50)
@@
$(\frac{K-\alpha}{\alpha})^{2}e^{-2u}\leq\epsilon.$ (51)
@@
$u\geq\frac{1}{2}\log\frac{1}{\epsilon}+\log\frac{K-\alpha}{\alpha}.$ (52)
@@
$\frac{1}{2}\log\frac{\eta t}{KM}+\log\frac{K-\alpha}{\alpha}\geq\frac{1}{2}\log\frac{1}{\epsilon}+\log\frac{K-\alpha}{\alpha},$ (53)
@@
${\tilde{{\bm{p}}_{i}^{n}}}^{\top}{\bm{W}}^{(t)}\tilde{{\bm{p}}}_{query}^{n}\geq\log\frac{1}{\epsilon}.$ (55)
@@
${\tilde{{\bm{p}}_{i}^{n}}}^{\top}{\bm{W}}^{(t)}\tilde{{\bm{p}}}_{query}^{n}\geq\Theta(1)\cdot\log\frac{1}{\epsilon}=\Theta(\log\frac{1}{\epsilon}).$ (56)
@@ E.2 Proof of Theorem 2
$\tilde{{\bm{p}}_{i}}=\begin{pmatrix}{\bm{\mu}}_{j}^{\prime}\\ {\bm{y}}_{i,1}\\ \end{pmatrix}+\begin{pmatrix}{\bm{\delta}}_{i}\\ \bm{0}\\ \end{pmatrix}+{\bm{c}}_{i\mod K}.$ (65)
@@
$=$ $(({{\bm{\mu}}_{j}^{\prime}}^{\top},{\bm{y}}_{i,1}^{\top})+{\bm{c}}_{i\mod K}^{\top}))({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})\tilde{{\bm{p}}}_{query}.$
Since that $\bm{\lambda}_{j}$ is orthogonal to all the ${\bm{\mu}}_{i}$, $i\in[M]$, we have similar conclusion for $\bm{\lambda}_{j}$ as ${\bm{\delta}}_{i}$, i.e.,
$(\bm{\lambda}_{j}^{\top},0^{\top}){\bm{W}}^{(0)}\tilde{{\bm{p}}_{i}}\lesssim\xi,$ (69)
and
$(\bm{\lambda}_{j}^{\top},0^{\top})\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}}\tilde{{\bm{p}}}_{query}=0.$ (70)
Let ${\bm{\mu}}_{j}^{\prime}=\bm{\lambda_{j}}+\tilde{{\bm{\mu}}}_{j}=\bm{\lambda_{j}}+\sum_{i=1}^{M^{\prime}}k_{j,i}{\bm{\mu}}_{i}$. Then, we have
@@
$=$ $((\bm{\lambda_{j}}^{\top}+\sum_{i=1}^{M^{\prime}}k_{j,i}{\bm{\mu}}_{i}^{\top},{\bm{y}}_{i,1}^{\top})+{\bm{c}}_{i\mod K}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})((\bm{\lambda_{j}}^{\top}$
$+\sum_{i=1}^{M^{\prime}}k_{j,i}{\bm{\mu}}_{i}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$
$=$ $\sum_{i=1}^{M^{\prime}}k_{j,i}^{2}(({\bm{\mu}}_{i}^{\top},{\bm{y}}_{i,1}^{\top})+{\bm{c}}_{i\mod K}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})(({\bm{\mu}}_{i}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$
$+\sum_{i\neq i^{\prime}}k_{j,i}k_{j,i^{\prime}}(({\bm{\mu}}_{i}^{\top},{\bm{y}}_{i,1}^{\top})+{\bm{c}}_{i\mod K}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})(({\bm{\mu}}_{i^{\prime}}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$
$\geq$ $C\cdot\Theta(\log\frac{1}{\epsilon})-\Theta(\xi)$
$=$ $\Theta(\log\frac{1}{\epsilon}),$
where the second to last step is by Theorem 1. The last step holds if $C\geq\Theta(\log^{-1}(1/\epsilon))$. Since the gradient updates for different TRR patterns are very close to each other, we have that $\sum_{i\neq i^{\prime}}|k_{j,i}k_{j,i^{\prime}}|\leq 1$ and
$\sum_{i\neq i^{\prime}}k_{j,i}k_{j,i^{\prime}}(({\bm{\mu}}_{i}^{\top},{\bm{y}}_{i,1}^{\top})+{\bm{c}}_{i\mod K}^{\top})({\bm{W}}^{(0)}+\frac{\eta}{B}\sum_{n\in\mathcal{B}_{b}}\sum_{b=0}^{T-1}\frac{\partial\ell(\Psi;{\bm{P}}^{n},{\bm{z}}^{n})}{\partial{\bm{W}}^{(b)}})(({\bm{\mu}}_{i^{\prime}}^{\top},\bm{0}^{\top})+{\bm{c}}_{1})^{\top}$ (72)
$\lesssim$ $\Theta(1)\cdot\frac{\tilde{{\bm{p}}_{s}}^{\top}{\bm{W}}^{(T)}\tilde{{\bm{p}}}_{query}}{\log\frac{1}{\epsilon}},$
where $\tilde{{\bm{p}}_{s}}$ shares the same TSR pattern and the same step as $\tilde{{\bm{p}}}_{query}$. Hence, for $\tilde{{\bm{p}}}_{i}$ that shares a different TSR pattern with $\tilde{{\bm{p}}}_{query}$,
$\tilde{{\bm{p}}_{i}}^{\top}{\bm{W}}^{(T)}\tilde{{\bm{p}}}_{query}\lesssim\Theta(1).$ (73)
@@
$\tilde{{\bm{p}}}_{j}^{\top}{\bm{W}}^{(T)}(({\bm{v}}_{k}^{\top},\bm{0}^{\top})^{\top}+{\bm{c}}_{k}^{\top})^{\top}=\Theta(\log\frac{1}{\epsilon}),$ (76)